OpenStudy (anonymous):

Please help! Evaluate without using trigonometric tables on calc: (sin45degrees)(cos30degrees)/sec45degrees a. 1/4 b. sqrt.3/4 c. sqrt.3/2 d. sqrt.3

OpenStudy (australopithecus):

sin(45)cos(30)/sec(45) sec(x) = 1/cos(x) so we have sin(45)cos(30)/1/cos(45)

OpenStudy (anonymous):

So b.?

OpenStudy (australopithecus):

sin(45) = (2)^(1/2)/2 cos(45) = 2^(1/2)/2 cos(30) = 3^(1/2)/2

OpenStudy (australopithecus):

based on memory

OpenStudy (australopithecus):

so you have \[(\frac{\sqrt{2}}{2})(\frac{\sqrt{2}}{2})(\frac{\sqrt{3}}{2})\]

OpenStudy (australopithecus):

as sin(45)cos(30)/1/cos(45) = sin(45)cos(30)cos(45)

OpenStudy (anonymous):

\[\large \sin(45) = \cos(45) = \frac{1}{\sqrt{2}}\] \[\cos(30) = \frac{\sqrt{3}}{2}\]

OpenStudy (australopithecus):

(2)^(1/2)*(2)^(1/2) = 4^(1/2) = 2

OpenStudy (anonymous):

Thanks guy! So the final answer is C. sqrt.3/2

OpenStudy (australopithecus):

i think you are messing up your fraction mulitplication

OpenStudy (australopithecus):

Remember